OpenStudy (anonymous):

when solving a system of equations by graphing how is the solution found? sketch a example

OpenStudy (anonymous):

By looking at the point of intersection.

OpenStudy (anonymous):

ex. 2x + 7y = 1 -3x – 4y = 5

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

put the points in and look at the point of intersection which is (-3,1)

OpenStudy (anonymous):

how can constants be used to model a real-world situation

OpenStudy (anonymous):

It can show when online daily news papers in the US make the same profit as printed news papers in the US.

OpenStudy (anonymous):

I don't get it. Can you give me another example

OpenStudy (anonymous):

It can show when the average viewing hours for cable T.V. and satellite T.V. will the same number of viewers.

OpenStudy (anonymous):

This is my last question. Solve the following of systems of equations by graphing, you must graph. Y = -2x - 1 y = -3/2x -3

OpenStudy (anonymous):

put the numbers into https://www.desmos.com/calculator

OpenStudy (anonymous):

I did how do I save a graph on there

OpenStudy (anonymous):

I got t thanks

OpenStudy (anonymous):

You can print it or on the left i think there's a save button.

OpenStudy (anonymous):

And you can save your graph on this one

OpenStudy (anonymous):

Mason owns a collection of hardback books in a series about American history. Each book has dimensions as shown below. There are 7 books in the series, stacked side by side. What is the volume of the entire collection? a) one hundred sixty-one cubic inches b) eight hundred forty cubic inches c) nine hundred twenty-four cubic inches d) nine hundred sixty cubic inches

OpenStudy (anonymous):

medal

OpenStudy (anonymous):